OpenStudy (anonymous):

DIVIDE (-3+5i)/(5-4i)

OpenStudy (anonymous):

\[\frac{ (-3+5i)(5+4i) }{ 5^{2}+4^{2} }\]

OpenStudy (anonymous):

\[\frac{ -15+13i+20i ^{2} }{ 41 }\]

OpenStudy (anonymous):

So far so good.

OpenStudy (anonymous):

im confused about what to do with the i^2

OpenStudy (anonymous):

is it i^2 = 1

OpenStudy (anonymous):

\(i^2=-1\)

OpenStudy (anonymous):

ah lol working..

OpenStudy (anonymous):

\[\frac{ 13i-35 }{ 41 }\]

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

and i cant do anything else, at this time

OpenStudy (anonymous):

math 143

OpenStudy (anonymous):

thanks for your time, sir

OpenStudy (anonymous):

You're welcome, I didn't have to do anything.

OpenStudy (anonymous):

:)